# Mastering NIM and Impartial Games with Weak Neural Networks: An AlphaZero-inspired Multi-Frame Approach

### v2

@@ 8 Conclusion
We proposed a circuit-complexity model for practical neural inference: constant-depth, polynomial-size networks with fixed-precision parameters. We showed that these models are simulable by $\mathrm{AC}^{0}$, which makes global parity-type computations (and hence NIM’s nim-sum) infeasible in the strong, worst-case sense.
This yields a clean explanation for persistent learning failures under single-frame training: strong mastery of NIM requires reliable access to global nim-sum information, which fixed-precision bounded-depth networks cannot represent, and polynomial-time bounded search does not remove this obstruction. On the positive side, we proved that a short history window exposes locally computable nimber differences, and that a player can restore the zero nim-sum invariant by matching the observed nimber difference between consecutive frames.
Our experiments support the restoration viewpoint in a controlled setting. In a supervised teacher–random restoration dataset for 20-heap, 4-bit NIM, the one-frame policy remains near chance, while the two-frame policy learns restoration via a characteristic two-stage curve (an intermediate plateau around 0.54 followed by near-perfect accuracy) and achieves a clear head-to-head advantage over the one-frame baseline, especially when combined with bounded per-move MCTS (Section 6.2). At the same time, more end-to-end teacher-guided MCTS training remains a challenging regime to bootstrap from terminal win/loss signals, highlighting the value of intermediate, transition-aligned learning signals for restoration behaviour.
More broadly, the $\mathrm{AC}^{0}$ perspective offers a diagnostic tool: when learning fails, the bottleneck may be a parity-like global constraint rather than a lack of data or optimization. Multi-frame inputs and locally checkable invariants provide a principled route to bypass such barriers in other combinatorial domains.
Appendix A Formal definitions of fixed-precision neural models
This appendix provides formal definitions of the fixed-precision neural models used throughout the paper.
A.1 Preliminaries: fixed precision and representable parameters
@@
$\mathcal{Q}_{W,D}=\Big\{\frac{k}{D}:k\in\{0,1,\dots,W\}\Big\}.$
In particular, the set of allowable weights/thresholds does not depend on input length $n$. This is the mathematical formalization of “32-bit precision” (or any constant word size): there are only finitely many representable constants.
@@ A.2 Feedforward model
Definition A.1 (Fixed-precision constant-depth neural network).
A fixed-precision $\mathsf{NN}$ of depth $L$ on $n$ Boolean inputs is specified by:
@@ A.3 Recurrent model (finite window)
Definition A.2 (Fixed-precision recurrent neural network with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-precision $\mathsf{RNN}$ consists of:
@@
  • fixed-precision parameters (weights/thresholds in $\mathcal{Q}_{W,D}$ and optional input negations) shared across time,
such that each update $h_{t+1}$ and $o_{t}$ is computed by a constant-depth fixed-precision threshold network over inputs drawn from $\{x_{t^{\prime}},h_{t^{\prime}}:t-T<t^{\prime}\leq t\}$. Equivalently, unrolling for $T$ steps yields a constant-depth feedforward computation with depth $O(L\cdot T)$.
@@ A.4 Attention-style model (finite window)
Definition A.3 (Fixed-precision LTST model with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-precision $\mathsf{LTST}$ model processes a sequence $(x_{t-T+1},\dots,x_{t})$ by:
  • computing short-term and long-term attention scores using fixed-precision parameters in $\mathcal{Q}_{W,D}$,
  • combining attended summaries through a constant number of layers of fixed-precision threshold computations,
@@
Appendix B An $\mathrm{AC}^{0}$ construction for simulating fixed-precision networks
@@
B.1 One threshold unit is in $\mathrm{AC}^{0}$ under fixed precision
We first show that a single fixed-precision threshold unit (as in Definition A.1) is computed by a depth-2 $\mathrm{AC}^{0}$ circuit of polynomial size.
Lemma B.1 (Fixed-precision threshold gate $\in\mathrm{AC}^{0}$).
@@ B.1 One threshold unit is in $\mathrm{AC}^{0}$ under fixed precision
Then $f$ is computable by a depth-2 $\mathrm{AC}^{0}$ circuit of size $O(m^{K})$, where $K:=WD$ is a constant (depending only on $W,D$).
@@
Because each $a_{j}$ is a nonnegative integer and each active input contributes at least $1$ to the left-hand side whenever $a_{j}\geq 1$, any satisfying assignment contains a subset of at most $b$ active variables whose coefficients already sum to at least $b$. More formally: if $\sum_{j}a_{j}z_{j}\geq b$, let $S=\{j:z_{j}=1\text{ and }a_{j}\geq 1\}$. Choose any multiset-free subset $T\subseteq S$ by greedily adding indices until the partial sum reaches $\geq b$. Since each added index increases the sum by at least $1$, this process stops after at most $b\leq W$ steps, hence $|T|\leq b\leq W\leq WD=K$.
@@ B.2 From one unit to a full network
Lemma B.3 (One layer of fixed-precision units is in $\mathrm{AC}^{0}$).
@@ Appendix C Implementation details (overview)
This appendix summarizes the components needed to implement the two-frame delta computation and the local restoration response within $\mathrm{AC}^{0}$ constraints.
@@
All subroutines are implementable in $\mathrm{AC}^{0}$ because comparisons and equality on $O(\log m)$-bit blocks are depth-2 with unbounded fan-in.

### v3

@@ 9 Conclusion
We model fixed-latency neural inference in a fixed-scale, bounded-range quantised regime and show that such models are simulable by $\mathrm{AC}^{0}$ (Theorem 2.4). For NIM this yields a sharp single-frame barrier: optimal play depends on the global nim-sum, a parity-type invariant outside $\mathrm{AC}^{0}$.
We then show how interaction can bypass this barrier. Rollouts amplify effective depth, but with a single policy head every rollout reveals the same linear functional of the nim-sum (Proposition 5.10), so increased search budget does not recover the missing bits. Multiple policy heads break this bottleneck by providing independent rollout channels (Proposition 5.9). Alternatively, giving the agent two consecutive frames exposes the nimber difference $\Delta(P,P^{\prime})$, enabling a restoration policy that maintains $\operatorname{nim}=0$ and acts as a universal verifier (Propositions 5.5 and 5.13).
Our experiments support the restoration viewpoint in a controlled setting. On 20-heap, 4-bit NIM with $10^{6}$ supervised restoration examples, a one-frame policy stays near chance while a two-frame policy reaches near-perfect restoration accuracy (Section 7.2). In a complementary single-frame evaluation, a multi-head FSM shootout protocol reaches perfect win/loss position classification while the single-head baseline remains bottlenecked by one-bit information (Section 7.3). At the same time, the difficulty of end-to-end MCTS training highlights an optimisation gap: representational sufficiency does not by itself guarantee that standard RL optimisers will discover the solution from sparse rewards.
Overall, circuit complexity offers a concrete design principle for reliability beyond games. In domains such as logical verification, compiler optimisation, or symbolic reasoning, optimal decisions often hinge on global invariants (e.g. satisfiability or cycle structure) that can be hard for fixed-scale quantised bounded-depth inference. Our results suggest that rather than only scaling depth/width/search, a more effective bypass is to restructure inputs so local transition signals are explicit, converting a “state solving” problem into a “transition verification” problem.
Appendix A Formal definitions of fixed-scale quantised neural models
This appendix provides formal definitions of the fixed-scale quantised neural models used throughout the paper.
A.1 Preliminaries: fixed-scale quantisation and representable parameters
@@
$\mathcal{Q}_{W,D}=\Big\{\frac{\ell}{D}:\ell\in\{0,1,\dots,W\}\Big\}.$
In particular, the set of allowable weights/thresholds does not depend on input length $n$. This models fixed-scale quantisation with bounded range: values lie on a fixed grid (step $1/D$) and saturate at magnitude $W/D$, independent of the input length $n$. This is closer to deployment-time fixed-point/INT8-style inference with fixed scaling factors than to floating-point arithmetic. In particular, mere finiteness of a floating-point format is not what drives our $\mathrm{AC}^{0}$ containment; the key is that the effective range/scale of thresholds does not grow with fan-in as $n$ increases.
@@ A.2 Feedforward model
Definition A.1 (Fixed-scale quantised constant-depth neural network).
A fixed-scale quantised $\mathsf{NN}$ of depth $L$ on $n$ Boolean inputs is specified by:
@@ A.3 Recurrent model (finite window)
Definition A.2 (Fixed-scale quantised recurrent neural network with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-scale quantised $\mathsf{RNN}$ consists of:
@@
  • fixed-scale quantised parameters (weights/thresholds in $\mathcal{Q}_{W,D}$ and optional input negations) shared across time,
such that each update $h_{t+1}$ and $o_{t}$ is computed by a constant-depth fixed-scale quantised threshold network over inputs drawn from $\{x_{t^{\prime}},h_{t^{\prime}}:t-T<t^{\prime}\leq t\}$. Equivalently, unrolling for $T$ steps yields a constant-depth feedforward computation with depth $O(L\cdot T)$.
@@ A.4 Attention-style model (finite window)
Definition A.3 (Fixed-scale quantised LTST model with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-scale quantised $\mathsf{LTST}$ model processes a sequence $(x_{t-T+1},\dots,x_{t})$ by:
  • computing short-term and long-term attention scores using fixed-scale quantised parameters in $\mathcal{Q}_{W,D}$,
  • combining attended summaries through a constant number of layers of fixed-scale quantised threshold computations,
@@
Appendix B An $\mathrm{AC}^{0}$ construction for simulating fixed-scale quantised networks
@@
B.1 One threshold unit is in $\mathrm{AC}^{0}$ under fixed-scale quantisation
We first show that a single fixed-scale quantised threshold unit (as in Definition A.1) is computed by a depth-2 $\mathrm{AC}^{0}$ circuit of polynomial size.
Lemma B.1 (Fixed-scale quantised threshold gate $\in\mathrm{AC}^{0}$).
@@ B.1 One threshold unit is in $\mathrm{AC}^{0}$ under fixed-scale quantisation
Then $f$ is computable by a depth-2 $\mathrm{AC}^{0}$ circuit of size $O(m^{K})$, where $K:=W$ is a constant (depending only on $W,D$).
@@
Because each $a_{j}$ is a nonnegative integer and each active input contributes at least $1$ to the left-hand side whenever $a_{j}\geq 1$, any satisfying assignment contains a subset of at most $b$ active variables whose coefficients already sum to at least $b$. More formally: if $\sum_{j}a_{j}z_{j}\geq b$, let $S=\{j:z_{j}=1\text{ and }a_{j}\geq 1\}$. Choose any multiset-free subset $T\subseteq S$ by greedily adding indices until the partial sum reaches $\geq b$. Since each added index increases the sum by at least $1$, this process stops after at most $b$ steps. Because $b\leq W$, we get $|T|\leq W=:K$.
@@ B.2 From one unit to a full network
Lemma B.3 (One layer of fixed-scale quantised units is in $\mathrm{AC}^{0}$).
@@ Appendix C Implementation details (overview)
This appendix summarises the components needed to implement the two-frame delta computation and the local restoration response within $\mathrm{AC}^{0}$ constraints.
@@
All subroutines are implementable in $\mathrm{AC}^{0}$ because comparisons and equality on $O(\log m)$-bit blocks have constant-depth implementations (e.g., depth-3 with unbounded fan-in).
@@ Appendix D Weak mastery examples
Appendix F Additional proofs for the rollout analysis
Proof of Proposition 5.9.
Fix a position $P$ with heap sizes $h_{1},\dots,h_{N}\in\{0,\dots,2^{B}-1\}$ and let $\sigma=\operatorname{nim}(P)=h_{1}\oplus\cdots\oplus h_{N}$. For each $j\in\{0,\dots,B-1\}$ define the deterministic rollout policy $\pi_{j}$:
“Subtract $2^{j}$ from the leftmost heap of size at least $2^{j}$; if none exists, subtract $1$ from the leftmost nonzero heap.”
Each $\pi_{j}$ is $\mathrm{AC}^{0}$-computable because it performs constant-depth comparisons to the threshold $2^{j}$ and selects the leftmost qualifying heap.
Consider a rollout under $\pi_{j}$ from $P$. While some heap is at least $2^{j}$, every move subtracts exactly $2^{j}$ from some heap, so the number of such moves is $L_{1}=\sum_{i=1}^{N}\lfloor h_{i}/2^{j}\rfloor$. After this phase all heaps satisfy $h_{i}<2^{j}$, and $\pi_{j}$ falls back to subtracting $1$ until termination, for a further $L_{2}=\sum_{i=1}^{N}(h_{i}\bmod 2^{j})$ moves. Thus the total number of moves is $L=L_{1}+L_{2}$, and (normal play) the first player wins iff $L$ is odd, i.e. iff $L\bmod 2=1$.
For each $i$, the parity of $\lfloor h_{i}/2^{j}\rfloor$ equals the $j$th bit of $h_{i}$, denoted $h_{i}^{(j)}$, so $L_{1}\bmod 2=\bigoplus_{i}h_{i}^{(j)}=\sigma_{j}$. Similarly, $(h_{i}\bmod 2^{j})\bmod 2=h_{i}^{(0)}$, so $L_{2}\bmod 2=\bigoplus_{i}h_{i}^{(0)}=\sigma_{0}$. Hence $\mathrm{RO}_{\pi_{0}}(P)=\sigma_{0}$ and $\mathrm{RO}_{\pi_{j}}(P)=\sigma_{j}\oplus\sigma_{0}$ for $j\geq 1$. From these $B$ outcomes the agent recovers $\sigma$.
To select a winning move, the agent enumerates all legal moves $m$ from $P$ (at most $N(2^{B}-1)$ possibilities), runs the $B$ rollouts from the successor position $P_{m}$, reconstructs $\sigma(P_{m})$, and accepts the first $m$ with $\sigma(P_{m})=0$. Each rollout has length $O(N2^{B})$, so the total computation is $O\bigl(N2^{B}\cdot B\cdot N2^{B}\bigr)=O(BN^{2}\cdot 4^{B})$. ∎
Appendix G Proof of Theorem 5.14
We prove each part of the theorem separately.
Proof of Theorem 5.14.
Part (i): Single-frame, single-policy limitation.
Let $\mathcal{A}=(E,\pi_{j},S)$ be a single-frame $\mathrm{AC}^{0}$ agent with value head $E$, a single rollout policy $\pi_{j}$, and search budget $S$. We show that $\mathcal{A}$ identifies a winning move on a uniformly random $N$-position with probability at most $2^{-(B-1)}+\mathrm{negl}(N)$.
Step 1: All rollout outcomes reduce to one unknown bit. By Proposition 5.10, the rollout outcome from any position $Q$ under $\pi_{j}$ satisfies $\mathrm{RO}_{\pi_{j}}(Q)=\ell_{j}(\sigma(Q))$, where $\ell_{j}\colon\mathbb{F}_{2}^{B}\to\mathbb{F}_{2}$ is a fixed $\mathbb{F}_{2}$-linear functional of the nim-sum $\sigma(Q)=\operatorname{nim}(Q)$.
During search the agent may test candidate moves $m$ from the root position $P$ by running rollouts from the successor positions $P_{m}$. Since $\sigma(P_{m})=\sigma(P)\oplus(h_{i}\oplus h_{i}^{\prime})$ for a move changing heap $i$ from $h_{i}$ to $h_{i}^{\prime}$, linearity gives
$\mathrm{RO}_{\pi_{j}}(P_{m})\;=\;\ell_{j}\bigl(\sigma(P)\bigr)\;\oplus\;\underbrace{\ell_{j}(h_{i}\oplus h_{i}^{\prime})}_{\text{known from move}~m},$
so every rollout outcome is a known affine function of the single unknown bit $\beta:=\ell_{j}(\sigma(P))$. Regardless of the budget $S$ and the choice of candidate moves, the agent learns at most the value of $\beta$.
Step 2: The agent’s total information. After all rollout queries the agent possesses:
  1. (a)
    the position $P$ (all $NB$ input bits),
  2. (b)
    the output $E(P)$ of the $\mathrm{AC}^{0}$ value head,
  3. (c)
    the single bit $\beta$.
Its decision function is therefore $f\colon\{0,1\}^{NB}\times\{0,1\}\to\{\text{moves}\}$. For each fixed $\beta\in\{0,1\}$, the map $P\mapsto f(P,\beta)$ is computable by an $\mathrm{AC}^{0}$ circuit (it composes $E$, the rollout policy, and Boolean operations on the known correction bits).
Step 3: One linear functional is insufficient for move selection. A winning move from an $N$-position $P$ requires finding a heap $i$ and setting $x_{i}^{\prime}:=x_{i}\oplus\sigma(P)$ with $x_{i}^{\prime}<x_{i}$. This depends on the exact $B$-bit value of $\sigma(P)\in\{1,\dots,2^{B}-1\}$. The bit $\beta=\ell_{j}(\sigma(P))$ imposes one $\mathbb{F}_{2}$-linear constraint, partitioning the nonzero nim-sum values into two fibres: one of size $2^{B-1}$ and one of size $2^{B-1}-1$ (since $\sigma=0$ belongs to exactly one fibre but is excluded for $N$-positions). Within a fibre, distinct values $\sigma\neq\sigma^{\prime}$ generically require different winning moves, because $x_{i}\oplus\sigma\neq x_{i}\oplus\sigma^{\prime}$ whenever $\sigma\neq\sigma^{\prime}$.
Step 4: Probability bound via $\mathrm{AC}^{0}$ correlation limits. Draw $P$ uniformly at random from $N$-positions with $N$ heaps of bit-width $B$. For large $N$ and fixed $B$, the conditional distribution of $\sigma(P)$ given $\beta$ is exponentially close (in $N$) to uniform over the relevant fibre. (This follows from standard concentration: the nim-sum is the XOR of $N$ independent $B$-bit values, and conditioning on one linear functional leaves the remaining B1B-1 bits nearly uniform.)
For each fixed $\beta$, the agent commits to a move via $f(P,\beta)$. Since $P\mapsto f(P,\beta)$ is $\mathrm{AC}^{0}$, the Linial–Mansour–Nisan theorem [8] and Håstad’s switching lemma [7] imply exponentially small correlation with any parity-type predicate on the $NB$ input bits. In particular, $f(\cdot,\beta)$ cannot distinguish among the $2^{B-1}-1$ possible nonzero nim-sum values consistent with $\beta$, except with advantage $\mathrm{negl}(N)$.
The probability of selecting a winning move is therefore at most
$\frac{1}{2^{B-1}-1}+\mathrm{negl}(N)\;\leq\;2^{-(B-1)}+\mathrm{negl}(N).$
Part (ii): Single-frame, multi-policy-head sufficiency.
This is Proposition 5.9, proved in full in Appendix F. The $B$ rollout heads $\pi_{0},\dots,\pi_{B-1}$ yield outcomes
$\sigma_{0}=\mathrm{RO}_{\pi_{0}}(P),\qquad\sigma_{j}=\mathrm{RO}_{\pi_{j}}(P)\oplus\sigma_{0}\quad(j\geq 1),$
from which the search controller recovers all $B$ bits of $\sigma(P)$. The agent then enumerates candidate moves $m$, runs the $B$ rollouts from each successor $P_{m}$, reconstructs $\sigma(P_{m})$, and accepts the first $m$ with $\sigma(P_{m})=0$. Total computation: $O(BN^{2}\cdot 4^{B})$.
Part (iii): Two-frame, single-policy sufficiency.
Let $\pi^{*}$ denote the two-frame restoration policy of Proposition 5.5, augmented with a fixed fallback rule for the opening move (where no previous-frame transition is available). By Proposition 5.13, $\pi^{*}$ is a universal verifier: $\mathrm{RO}_{\pi^{*}}(Q)=0$ iff $\operatorname{nim}(Q)=0$.
The agent proceeds from an $N$-position $P$ as follows.
  1. (i)
    Find a $P$-position successor. Enumerate all legal moves $m$ from $P$ (at most $N(2^{B}-1)$ candidates). For each, run a single rollout under $\pi^{*}$ from the successor position $P_{m}$. Accept the first move $m$ for which $\mathrm{RO}_{\pi^{*}}(P_{m})=0$, indicating $\operatorname{nim}(P_{m})=0$. Since $P$ is an $N$-position, Theorem 4.4 guarantees at least one such move exists, so the search terminates successfully.
  2. (ii)
    Maintain $\operatorname{nim}=0$ by restoration. After the initial move the opponent faces a $P$-position $P_{m}$. Whatever the opponent plays, the agent observes the two-frame transition $(P_{m},\,P_{m}^{\prime})$ and applies the restoration rule of Proposition 5.5: it computes $\delta=\Delta(P_{m},P_{m}^{\prime})$ and responds with a move that realises the same nimber difference $\delta$, returning to $\operatorname{nim}=0$. No further rollouts are needed; all subsequent moves use only the $\mathrm{AC}^{0}$-computable two-frame restoration.
The computation cost is dominated by the initial search: at most $N(2^{B}-1)$ candidate moves, each tested with one rollout of length at most $O(N\cdot 2^{B})$ steps, giving total computation $O(N^{2}\cdot 4^{B})$. ∎